# Discrete flavour and CP symmetries in light of JUNO and neutrino global fit

### v1

@@ 6 Summary and conclusion
In order to further quantitatively assess how well a viable breaking pattern can describe the current experiment data (NuFIT and JUNO 59.1-day results), we perform a parameter scan and compute the minimum $\chi^{2}$ (defined in Eq. (5.1)) value for those one- and two-parameter mixing patterns. This analysis is restricted to mixing patterns arising from $\Delta(6n^{2})\rtimes H_{CP}$ with $n\leq 4$ and from $D_{n}\rtimes H_{CP}$ with $n\leq 8$. The best-fit values of input parameter and mixing parameters for mixing patterns with one input parameter for both NO and IO cases are shown in table 4. The best-fit values of input parameters and mixing parameters for mixing patterns with two input parameters arise from $A_{5}\rtimes H_{CP}$, $\Sigma(168)\rtimes H_{CP}$, $\Delta(6n^{2})\rtimes H_{CP}$ and $D_{n}\rtimes H_{CP}$ are shown in tables 5, 6 and 7. We scanned the parameter spaces of each mixing matrix to identify viable ranges consistent with neutrino oscillation data at the $3\sigma$ level. This procedure yielded predictions for lepton mixing parameters with the results detailed in tables 9, 10, 11 and 12. Furthermore, we can pin down the minimal value of $\chi^{2}_{\text{min}}$ for each mixing matrix relevant to the mixing angles and the Dirac CPV phase. The results for mixing patterns with one and two free parameters are shown in figures 1 and 3, respectively. We then assessed the robustness of these predictions under future experimental conditions. This assessment assumed that: (1) the current best-fit values for the mixing angles remain unchanged, and (2) the relative $1\sigma$ uncertainties on $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ and $\delta_{CP}$ are scaled to $0.5\%$, $3\%$ and $12^{\circ}$ of their current best-fit values, respectively. These projected uncertainties, corresponding to 6 years of JUNO [31] data and 15 years of DUNE [92] and T2HK [93] data, are visualized in figures 2, 4, 5, 6 and 7 in which the likelihood profiles for $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ and $\delta_{CP}$ are displayed for selected viable patterns with $\chi^{2}_{\min}\leq 9$.
@@
For the one-parameter mixing patterns, all five candidates are capable of accommodating the current NuFIT data and yield sharply constrained predictions for the lepton mixing parameters in both the NO and IO neutrino mass spectra. The only exceptions are $U^{IV}_{1}$ and $U^{IV}_{2}$, whose predictions imply a CP-conserving Dirac phase $\delta_{CP}$; these remain compatible with the NuFIT constraints only in the NO case. The smallest $\Delta(6n^{2})$ group that can realize the viable patterns $U^{I}$ and $U^{II}$ is $\Delta(24)\cong S_{4}$ with $n=2$. We note that the minimal group capable of generating $U^{II}$ is $A_{4}$ if the nontrivial singlet representations are absent from the model [61]. For $n=2$, the choice $(\varphi_{1},\varphi_{2})=\left(-\frac{\pi}{2},\,\frac{\pi}{2}\right)$ reduces $U^{I}$ to the well-known TM1 mixing form. The recent JUNO 59.1-day results further sharpen these constraints. At the $3\sigma$ level, the viable parameter space of one-parameter patterns is significantly reduced. In the NO case, only TM1 mixing ($U^{I}$) and the patterns $U^{IV}_{1}$ ($U^{IV}_{2}$), which feature fixed columns $(\phi_{g},\phi_{g}-1,1)^{T}/2$ ($(\phi_{g},1,\phi_{g}-1)^{T}/2$), remain compatible after incorporating the JUNO likelihood, as shown in figure 2. For the IO spectrum, $U^{I}$ (TM1) is the sole surviving option. Mixing patterns such as $U^{II}$ (TM2), $U^{III}$ (GR2), and $U^{V}$ (TM2) are disfavored once the current JUNO constraint on $\sin^{2}\theta_{12}$ is taken into account. We note that, for the mixing pattern $U^{I}$, higher values of $n$ in $\Delta(6n^{2})$ admit additional viable textures—beyond the TM1 form—that are consistent with current JUNO data. Looking into the future, the full set of one-parameter patterns is expected to come under decisive scrutiny: the projected precision of JUNO on $\sin^{2}\theta_{12}$, together with forthcoming measurements of $\sin^{2}\theta_{23}$ and $\delta_{CP}$ from DUNE and T2HK, will provide a definitive test of these scenarios.
@@
Nearly all two-parameter neutrino mixing patterns remain consistent with current NuFIT data and JUNO 59.1-day measurement of $\sin^{2}\theta_{12}$. Moreover, future high-precision JUNO results are unlikely to exclude the majority of these two-parameter scenarios. Distinguishing among these patterns will require improved long-baseline measurements of $\sin^{2}\theta_{23}$ and $\delta_{CP}$. For mixing patterns from $A_{5}\rtimes H_{CP}$, those with fixed elements $\frac{\phi_{g}}{2}$, $\frac{1}{2}$ and $\frac{1}{2\phi_{g}}$ are viable. The best experimental agreement is furnished by the $U^{VI}_{6,1}$ pattern, which has $\frac{1}{2\phi_{g}}$ fixed at the $(21)$ entry and gives the best fit for NO data, but conserves the CP-symmetry. In contrast, $U^{VI}_{1}$, with the fixed element $\frac{\phi_{g}}{2}$ at the $(11)$ entry, offers a good fit to both NO and IO and predicts nontrivial CP violation. For the mixing patterns derived from $\Sigma(168)\rtimes H_{CP}$, the fixed entry is $\frac{1}{2}$. The smallest $\chi^{2}_{\rm min}$ among permutations occurs for $U^{VII}_{2,2}$, with the fixed $(31)$ entry. Among the three mixing patterns $U^{VIII}$, $U^{IX}$ and $U^{X}$ obtained from $\Delta(6n^{2})\rtimes H_{CP}$—with fixed entries $\cos\varphi_{1}$, $\frac{1}{2}$ and $\frac{1}{\sqrt{2}}$, respectively—the minimal flavour groups realizing these patterns with conserved CP phases are $\Delta(54)$ for $U^{VIII}$ and $S_{4}$ for both $U^{IX}$ and $U^{X}$. Allowing for CP violation requires extending the symmetry to $\Delta(96)$, and then $U^{VIII}_{9}$ for $\varphi_{1}=\varphi_{2}=\frac{\pi}{4}$ and $U^{X}_{2}$ for $(\varphi_{1},\varphi_{2})=(\frac{\pi}{4},0)$ give rise to $\frac{1}{\sqrt{2}}$ at $(33)$ entry, yielding the smallest $\chi^{2}_{\rm min}$. It is remarkable that the dihedral groups $D_{3}$ and $D_{4}$ can also give rise to the viable mixing pattern $U^{VIII}$ with CP-conserving and CP-violating phases, respectively.
@@
In summary, we have shown that the discrete flavour symmetries $A_{5}\rtimes H_{CP}$, $\Sigma(168)\rtimes H_{CP}$, $\Delta(6n^{2})\rtimes H_{CP}$ and $D_{n}\rtimes H_{CP}$, when broken to suitable residual symmetries in the charged lepton and neutrino sectors, give rise to a well defined set of viable neutrino mixing scenarios characterized by one or two real parameters. These constructions lead to sharp, testable predictions for the mixing angles $\theta_{12}$, $\theta_{23}$ and the Dirac phase $\delta_{CP}$. We find that the smallest flavour group capable of generating phenomenologically viable one-parameter mixing patterns is $S_{4}$ (of order 24). For two-parameter mixing patterns, the minimal realizations arise from the dihedral groups $D_{4}$ (order 8) or $D_{3}$ (order 6), the latter giving rise to CP-conserving Dirac and Majorana phases. The recent JUNO 59.1-day dataset has reduced the $3\sigma$ region of $\sin^{2}\theta_{12}$ with respect to that of NuFIT by a factor of roughly 1.3, exerting a particularly strong impact on one-parameter scenarios, while the constraints on two-parameter patterns remain comparatively mild. A decisive discrimination among the surviving mixing patterns will require the next generation of precision measurements. In particular, the full JUNO dataset, together with the upcoming long-baseline experiments DUNE and T2HK, will be essential for thoroughly testing and potentially ruling out, or providing evidence for, the residual symmetry-based frameworks, originating form the non-Abelian discrete symmetry approach to the lepton mixing problem in particle physics.
@@ Acknowledgements
CCL is supported by Natural Science Basic Research Program of Shaanxi (Program No. 2024JC-YBQN-0004), and the National Natural Science Foundation of China under Grant No. 12247103. JNL is supported by the Grant No. NSFC-12505133. GJD is supported by the National Natural Science Foundation of China under Grant Nos. 12375104 and Guizhou Provincial Major Scientific and Technological Program XKBF (2025)010. The work of S. T. P. was supported in part by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Staff Exchange grant agreement No 101086085 – ASYMMETRY, by the Italian INFN program on Theoretical Astroparticle Physics and by the World Premier International Research Center Initiative (WPI Initiative, MEXT), Japan.
@@
Appendix A Group theory of $A_{5}$, $\Sigma(168)$ and $\Delta(6n^{2})$
@@ Appendix B Allowed ranges of lepton mixing observables for viable one- and two-parameter patterns
NO $U^{I}$ $(-\frac{\pi}{2},\frac{\pi}{2})$ $[0.3171,0.3195]$ $[0.02030,0.02381]$ $0.5$ $1.5$ 0 0
$U^{II}_{1}$ $(0,\pi)$ $[0.3403,0.3415]$ $[0.02033,0.02379]$ $0.5$ $1.5$ 0 0
$U^{II}_{2}$ $(0,\pi)$ $[0.3402,0.3415]$ $[0.02030,0.02379]$ $0.5$ $1.5$ 0 0
$U^{III}$ $[0.2821,0.2831]$ $[0.02034,0.02386]$ $0.5$ $1.5$ 0 0
$U^{IV}_{1}$ $[0.3295,0.3319]$ $[0.02031,0.02383]$ $[0.515,0.532]$ 11 0 0
$U^{IV}_{2}$ $[0.3295,0.3319]$ $[0.02031,0.02383]$ $[0.468,0.485]$ 0 0 0
$U^{V}_{1}$ $[0.3407,0.3415]$ $[0.02153,0.02385]$ $[0.541,0.585]$ $1.210,1.382]^{*} [$$0.153,0.165]^{*} [$$0.544,0.565]^{*}$
$U^{V}_{2}$ $[0.3407,0.3415]$ $[0.02157,0.02385]$ $[0.435,0.459]$ $[1.618,1.710]$ $[0.835,0.841]$ $[0.435,0.446]$
IO $U^{I}$ $(-\frac{\pi}{2},\frac{\pi}{2})$ $[0.3169,0.3193]$ $[0.02062,0.02408]$ $0.5$ $1.5$ 0 0

### v2

@@ 6 Summary and conclusion
In order to further quantitatively assess how well a mixing determined by a given breaking pattern can describe the current experiment data (NuFIT and JUNO 59.1-day results), we perform a parameter scan and compute the minimum $\chi^{2}$ (defined in Eq. (5.1)) value for the derived one- and two-parameter mixing patterns. In the cases of $\Delta(6n^{2})\rtimes H_{CP}$ and $D_{n}\rtimes H_{CP}$, this analysis is restricted to groups respectively with $n\leq 4$ and $n\leq 8$. The best-fit values of the input parameter and the lepton mixing parameters for the mixing patterns with one input parameter are shown for both NO and IO cases in table 4. Those for the mixing patterns with two input parameters arising from the groups from $A_{5}\rtimes H_{CP}$, $\Sigma(168)\rtimes H_{CP}$, $\Delta(6n^{2})\rtimes H_{CP}$ and $D_{n}\rtimes H_{CP}$ are shown respectively in tables 5, 6 and 7.
We scanned the parameter spaces of each mixing matrix to identify viable ranges consistent with neutrino oscillation data at the $3\sigma$ level. This procedure yielded predictions for the lepton mixing parameters shown in detail in tables 9, 10, 11 and 12. Furthermore, we can determine the minimal value of $\chi^{2}_{\text{min}}$ for the mixing angles and the Dirac CPV phase for each considered mixing matrix. The results for the mixing patterns with one and two free parameters are shown in figures 1 and 3, respectively. We then assessed the robustness of these predictions under future projected experimental conditions. This assessment assumed that: (1) the current best-fit values for the mixing angles remain unchanged, and (2) the relative $1\sigma$ uncertainties on $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ and $\delta_{CP}$ are scaled to $0.5\%$, $3\%$ and $12^{\circ}$ of their current best-fit values, respectively. These projected uncertainties, corresponding to 6 years of JUNO [34] data and 15 years of DUNE [95] and T2HK [96] data, are visualized in figures 2, 4, 5, 6 and 7 in which the likelihood profiles for $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ and $\delta_{CP}$ are displayed for selected viable patterns with $\chi^{2}_{\min}\leq 9$.
@@
For the one-parameter mixing patterns, all five candidates are capable of accommodating the current NuFIT data and yield sharply constrained predictions for the lepton mixing parameters for both the NO and IO neutrino mass spectra. The only exceptions are $U^{IV}_{1}$ and $U^{IV}_{2}$, whose predictions imply a CP-conserving Dirac phase $\delta_{CP}$; this is compatible with the NuFIT constraints only in the NO case. The smallest $\Delta(6n^{2})$ group that can realize the viable patterns $U^{I}$ and $U^{II}$ is $\Delta(24)\cong S_{4}$ with $n=2$. We note that the minimal group capable of generating $U^{II}$ is $A_{4}$ if the nontrivial singlet representations are absent from the model [64]. For $n=2$, the choice of the relevant discrete parameters $(\varphi_{1},\varphi_{2})=\left(-\frac{\pi}{2},\,\frac{\pi}{2}\right)$ reduces $U^{I}$ to the well-known TM1 mixing form. The recent JUNO 59.1-day results further sharpen these constraints. At the $3\sigma$ level, the viable parameter space of the one-parameter patterns is significantly reduced. In the NO case, only TM1 mixing ($U^{I}$) and the patterns $U^{IV}_{1}$ ($U^{IV}_{2}$), which feature fixed columns $(\phi_{g},\phi_{g}-1,1)^{T}/2$ ($(\phi_{g},1,\phi_{g}-1)^{T}/2$), remain compatible after incorporating the JUNO likelihood, as shown in figure 2. For the IO spectrum, $U^{I}$ (TM1) is the sole surviving option. Mixing patterns such as $U^{II}$ (TM2), $U^{III}$ (GR2), and $U^{V}$ (TM2) are disfavored once the current JUNO constraint on $\sin^{2}\theta_{12}$ is taken into account. We note that, for the mixing pattern $U^{I}$, higher values of $n$ in $\Delta(6n^{2})$ admit additional viable textures—beyond the TM1 form—that are consistent with current JUNO data. Looking into the future, the full set of one-parameter patterns is expected to come under decisive scrutiny: the projected precision of JUNO on $\sin^{2}\theta_{12}$, together with forthcoming measurements of $\sin^{2}\theta_{23}$ and $\delta_{CP}$ from DUNE and T2HK, will provide a definitive test of these scenarios.
@@
Nearly all two-parameter neutrino mixing patterns remain consistent with current NuFIT data and JUNO 59.1-day measurement of $\sin^{2}\theta_{12}$. Moreover, future high-precision JUNO results are unlikely to exclude the majority of these two-parameter scenarios. Distinguishing among these patterns will require improved long-baseline measurements of $\sin^{2}\theta_{23}$ and $\delta_{CP}$. For the mixing patterns from $A_{5}\rtimes H_{CP}$, those with fixed elements $\frac{\phi_{g}}{2}$, $\frac{1}{2}$ and $\frac{1}{2\phi_{g}}$ are viable. The best experimental agreement is furnished by the $U^{VI}_{6,1}$ pattern, which has $\frac{1}{2\phi_{g}}$ fixed at the $(21)$ entry and gives the best fit for the NO data, while conserving the CP-symmetry. In contrast, $U^{VI}_{1}$, with the fixed element $\frac{\phi_{g}}{2}$ at the $(11)$ entry, offers a good fit in both NO and IO cases and predicts nontrivial CP violation effects in neutrino oscillations. For the mixing patterns derived from $\Sigma(168)\rtimes H_{CP}$, the fixed entry is $\frac{1}{2}$. The smallest $\chi^{2}_{\rm min}$ among permutations occurs for $U^{VII}_{2,2}$, with the fixed $(31)$ entry. Among the three mixing patterns $U^{VIII}$, $U^{IX}$ and $U^{X}$ obtained from $\Delta(6n^{2})\rtimes H_{CP}$—with fixed entries $\cos\varphi_{1}$, $\frac{1}{2}$ and $\frac{1}{\sqrt{2}}$, respectively—the minimal flavour groups realizing these patterns with CP-conserving Dirac and Majorana phases are $\Delta(54)$ for $U^{VIII}$ and $S_{4}$ for both $U^{IX}$ and $U^{X}$. Allowing for CP violation requires extending the symmetry to $\Delta(96)$. Then $U^{VIII}_{9}$ for $\varphi_{1}=\varphi_{2}=\frac{\pi}{4}$ and $U^{X}_{2}$ for $(\varphi_{1},\varphi_{2})=(\frac{\pi}{4},0)$ give rise to $\frac{1}{\sqrt{2}}$ at $(33)$ entry, yielding the smallest $\chi^{2}_{\rm min}$. It is remarkable that the dihedral groups $D_{3}$ and $D_{4}$ can also give rise to the viable mixing pattern $U^{VIII}$ with CP-conserving and CP-violating phases, respectively.
@@
In summary, we have shown that the discrete flavour symmetries $A_{5}\rtimes H_{CP}$, $\Sigma(168)\rtimes H_{CP}$, $\Delta(6n^{2})\rtimes H_{CP}$ and $D_{n}\rtimes H_{CP}$, when broken to suitable residual symmetries in the charged lepton and neutrino sectors, give rise to a well defined set of viable neutrino mixing scenarios characterized by one or two real parameters. These constructions lead to sharp, testable predictions for the mixing angles $\theta_{12}$, $\theta_{23}$ and the Dirac phase $\delta_{CP}$. We find that the smallest flavour group capable of generating phenomenologically viable one-parameter mixing patterns is $S_{4}$ (of order 24). For two-parameter mixing patterns, the minimal realizations arise from the dihedral groups $D_{4}$ (order 8) or $D_{3}$ (order 6), the latter giving rise to CP-conserving Dirac and Majorana phases. The recent JUNO 59.1-day dataset has reduced the $3\sigma$ allowed region of $\sin^{2}\theta_{12}$ with respect to that of NuFIT by a factor of roughly 1.3, exerting a particularly strong impact on one-parameter scenarios, while the constraints on two-parameter patterns remain comparatively mild. A decisive assessment of the phenomenological viability of the surviving mixing patterns will require the next generation of precision measurements. In particular, the full JUNO dataset, together with the upcoming long-baseline experiments DUNE and T2HK, will be essential for thoroughly testing and potentially ruling out, or providing evidence for, the residual symmetry-based frameworks, originating form the non-Abelian discrete symmetry approach to the lepton mixing problem in particle physics.
@@ Acknowledgements
CCL is supported by Natural Science Basic Research Program of Shaanxi (Program No. 2024JC-YBQN-0004), and the National Natural Science Foundation of China under Grant No. 12247103. JNL is supported by the Grant No. NSFC-12505133. GJD is supported by the National Natural Science Foundation of China under Grant Nos. 12375104, 12547106 and Guizhou Provincial Major Scientific and Technological Program XKBF (2025)010. The work of S. T. P. was supported in part by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Staff Exchange grant agreement No 101086085 – ASYMMETRY, by the Italian INFN program on Theoretical Astroparticle Physics and by the World Premier International Research Center Initiative (WPI Initiative, MEXT), Japan.
@@
Appendix A Group theory of $A_{5}$, $\Sigma(168)$, $\Delta(6n^{2})$ and $D_{n}$
@@ Appendix B Allowed ranges of lepton mixing observables for viable one- and two-parameter patterns
NO $U^{I}$ $(-\frac{\pi}{2},\frac{\pi}{2})$ $[0.3171,0.3195]$ $[0.02030,0.02388]$ $0.5$ $1.5$ 0 0
$U^{II}_{1}$ $(0,\pi)$ $[0.3402,0.3415]$ $[0.02030,0.02388]$ $0.5$ $1.5$ 0 0
$U^{II}_{2}$ $(0,\pi)$ $[0.3402,0.3415]$ $[0.02030,0.02388]$ $0.5$ $1.5$ 0 0
$U^{III}$ $[0.2821,0.2832]$ $[0.02030,0.02388]$ $0.5$ $1.5$ 0 0
$U^{IV}_{1}$ $[0.3295,0.3319]$ $[0.02030,0.02388]$ $[0.515,0.532]$ 11 0 0
$U^{IV}_{2}$ $[0.3295,0.3319]$ $[0.02030,0.02388]$ $[0.468,0.485]$ 0 0 0
$U^{V}_{1}$ $[0.3407,0.3415]$ $[0.02153,0.02388]$ $[0.541,0.585]$ $1.210,1.382]^{*} [$$0.153,0.165]^{*} [$$0.544,0.565]^{*}$
$U^{V}_{2}$ $[0.3407,0.3415]$ $[0.02153,0.02388]$ $[0.435,0.459]$ $[1.618,1.710]$ $[0.835,0.841]$ $[0.435,0.446]$
IO $U^{I}$ $(-\frac{\pi}{2},\frac{\pi}{2})$ $[0.3169,0.3193]$ $[0.02060,0.02409]$ $0.5$ $1.5$ 0 0